Nevyaya20:

The table and the graph each show a different relationship between the same two variables, x and y: A table with two columns and 5 rows is shown. The column head for the left column is x, and the column head for the right column is y. The row entries in the table are 3,180 and 4,240 and 5,300 and 6,360. On the right of this table is a graph. The x axis values are from 0 to 10 in increments of 2 for each grid line. The y axis values on the graph are from 0 to 350 in increments of 70 for each grid line. A line passing through the ordered pairs 2, 70 and 4, 140 and 6, 210 and 8, 280 is drawn. How much more would the value of y be in the table than its value on the graph when x = 11? (5 points) a 110 b 150 c 215 d 275

Floo:

\[\frac{ 240-180 }{ 4-3 }=60 \] point slope form = \[y-y _{1}\]\[=m(x-x _{1})\] y-180=60(x-3) y-180=60x-180 (slope above) y=60x y=60x11=660 the value of 660 at x=11 you do the rest